on Monday,Anna read 25% of her book. On Tuesday, she read 30% of her book. On Wednesday,she read 25% of her book.On Thursday,she read 25% of her book. She still has 10% of her book left to read.This is possible. expalin why
1
Expert's answer
2019-05-01T10:17:50-0400

Let x be the amount of book pages. If book is always the same size, Anna will read 25%+30%+25%+25%+10%=125%. So, the book must be reduced and percents for the rest part were described each day.

On Monday x*25%/100%=0.25x was read, rest: x-0.25x=0.75x;

On Tuesday, 30% of the rest was read:  0.75x*0.3=0.225x, rest: 0.75x-0.225x=0.525x

On Wednesday 25% of the rest was read: 0.525x*0.25=0.1325x, rest: 0.525x-0.1325x=0.3975x

On Thursday 25% of the rest was read: 0.25*0.3975x=0.098438x, rest: 0.3975x-0.098438x=0.3x

Therefore, it is possible that left 30%

Answer: It is possible if the book is reducible (each time the number of pages left in the book is reduced).